OpenStudy (anonymous):

Integral of x/sqrt (x^2+4)

OpenStudy (anonymous):

try \(u=x^2+4\)

OpenStudy (anonymous):

it's for trigonometric sustitution...

OpenStudy (anonymous):

lol why?

OpenStudy (anonymous):

you can do it in your head \[u=x^2+4, du =2xdx, \frac{1}{2}du=dx\] \[\frac{1}{2}\int \frac{du}{\sqrt{u}}=\sqrt{u}=\sqrt{x^2+4}\]

OpenStudy (anonymous):

we can do it with a trig sub if you like put \(x=\tan(\theta), dx=\sec^2(\theta)d\theta\)

OpenStudy (anonymous):

oops, forgot about the 4 put \[x=2\tan(\theta), dx=2\sec^2(\theta)d\theta\]

OpenStudy (anonymous):

good so far?

OpenStudy (anonymous):

yes, i did that, but i don't know what to do next \[\int\limits2\tan \theta /\sqrt{2\tan^2\theta+4}\]

OpenStudy (anonymous):

you forgot the \(2\sec^2(\theta)\) up top

OpenStudy (anonymous):

also you forgot to square the 2 inside the radical

OpenStudy (anonymous):

oh yes, i forgot that, anyway what should i do next?

OpenStudy (anonymous):

\[\int \frac{2\tan(\theta)2\sec^2(\theta)}{\sqrt{4\tan^2(\theta)+4}}d\theta\]

OpenStudy (anonymous):

then \(\sqrt{4\tan^2(\theta)+4}=\sqrt{4(\tan^2(\theta)+1)}=2\sec(\theta)\)

OpenStudy (anonymous):

cancel a 2 cancel a secant, get \[2\int \sec(\theta)\tan(\theta)d\theta\]

OpenStudy (anonymous):

a function whose anti derivative you know right?

OpenStudy (anonymous):

yes, thanks i appreciate it :)

OpenStudy (anonymous):

this is really a long chasing the tail method of doing it you are converting to trig and converting back, but no matter, if this is what you have to do, then you can do it

OpenStudy (anonymous):

yw